OpenStudy (anonymous):

Write an equation of the line that is parallel to the given line y = 2x + 3 and the given point (-4,1).

OpenStudy (anonymous):

A parallel line has the same slope so if you put everything into point slope form, you get: (y-1)=2(x+4)

OpenStudy (anonymous):

is that the answer because i have no idea on how to do this

OpenStudy (hba):

@ArkGoLucky She does not know how to find the slope even if you provide her with the formula.

OpenStudy (hba):

She is just looking for an answer.

OpenStudy (anonymous):

She gave me the slope

OpenStudy (anonymous):

sorry i dont know how to do this

OpenStudy (anonymous):

my answer is correct. You just have to simplify it to get it into point slope form(y=mx+b)

OpenStudy (anonymous):

ok thanks